# Useful entanglement can be extracted from noisy graph states

### v1

@@ Appendix A Post-measurement correlations and stabilizers compatible with the measurement scheme
Given a graph state |$\ket{G}$ of $N$ qubits and a certain measurement scheme defined in Lemma 1 by a set of qubits $I$ and operators $o_{i}$, then the subset of the stabilizer of |$\ket{G}$ compatible with the measurement pattern has $N-\lvert I\rvert$ generators, and therefore fully defines the post-measurement state.
@@
The stabilizer of |$\ket{G}$, up to the structure of the signs, can be viewed as a linear subspace of $\mathbb{Z}_{2}^{2N}$ – this is known as the binary representation of the stabilizer [5]. The identification is following: if the stabilizer operator is defined as $\prod_{j\in J}g_{j}$, the corresponding element of $\mathbb{Z}_{2}^{2N}$ is $(x_{1},\ldots,x_{N})\oplus(z_{1},\ldots,z_{N})$, with $x_{i}$ being 1 if and only if $i\in J$ and $\vec{z}=A_{G}\vec{x}\mod 2$, where $A_{G}$ is the adjacency matrix of $G$. Thus, if a stabilizer contains $X_{i}$, the corresponding numbers are $(x_{i},z_{i})=(1,0)$, for $Z_{i}$ it is $(x_{i},z_{i})=(0,1)$, and $Y_{i}$ corresponds to $(x_{i},z_{i})=(1,1)$; no operator at site $i$ is denoted by $(x_{i},z_{i})=(0,0)$. In this way, the additive algebra of $\mathbb{Z}_{2}^{2}$ mimics the product rules of Pauli operators with only the missing sign structure.
@@
The stabilizer of |$\ket{G}$ corresponds to a special $N$-dimensional subspace in $\mathbb{Z}_{2}^{2N}$ [5]. For any stabilizer operator to be compatible with the measurement scheme in the sense of Lemma 1, each operator of its Pauli string form with support in $I$ must be $o_{i}$: this introduces constraints to the set of compatible operators. However, these constraints translated into the language of $\mathbb{Z}_{2}^{2N}$ are linear: if $X_{i}$ is measured at site $i$, the additional constraint is $z_{i}=0$ (this encompasses the cases of $X_{i}$ and identity at site $i$), if $Z_{i}$ is measured, the constraint is $x_{i}=0$, and for $Y_{i}$ it is $x_{i}+z_{i}=0\mod 2$. All of them are independent, and there are $\lvert I\rvert$ of them, one for each measured site. Thus, the linear subspace of $\mathbb{Z}_{2}^{2N}$ corresponding to measurement-compatible stabilizers has dimension $N-\lvert I\rvert$; any basis of this subspace corresponds to generators of the post-measurement state stabilizer. ∎
@@
A.1 Patterns for specific graph families
@@
Both cases are consistent with a maximally entangled two-qubit state being produced in the terminal qubits: for the odd-$n$ case, the existence of stabilizers described by Equation 30 ensures that further measurement of $X$, $Y$ or $Z$ at qubit 00 is perfectly (anti)correlated with the outcome of measurement of the same operator at qubit $n+1$.
@@
This class of graphs includes the path graph (each layer has only a single qubit), the crazy graph (the layers 00 and $n+1$ contain one qubit each, every other has two) and twisted graphs (the qubit count alternates between 1 and 2) as well. Suppose that all the internal qubits in every layer are measured in the $X$ basis. The stabilizer operators that determine the remaining correlations must be consistent with this structure. A particularly simple set of such operators exists: first, observe that if a layer $k$ contains at least two qubits, $X_{k,i}X_{k,i^{\prime}}$ is a stabilizer operator. Thus, the signs of outcomes have to be equal within a layer. Therefore, for any choice of inner-layer qubit index $(i_{k})$ the direct analogue of Equation 28 holds:
@@
are stabilizer operators. For symmetry, we assume the twisted graph structure exists only for an odd number of layers, so that the terminal ones consist only of one qubit each. Thus, the terminal layers 00 and $n+1$ always consist only of one qubit each, and the inner indices are omitted. Consequently, if $s_{k}$ in Equation 30 are interpreted as signs of outcomes at layer $k$ (which are all equal within a layer due to the structure of the stabilizer operators mentioned previously), the same post-measurement stabilizer structure appears.
In the case of the crazy graph, an additional observation can be made: loss of any single internal (not adjacent to the terminal qubits) edge yields the same post-measurement stabilizer structure, since the indices $(i_{k})$ in Equation 32 can be chosen arbitrarily. Thus, the crazy graph family is especially tolerant to single edge loss: it does not affect the post-measurement state, if the outcomes are postselected on the ideal sign structure. This can be generalized to the loss of any subset of edges not adjacent to each other and the terminal qubits.
Appendix B Derivation for the noise models
B.1 Uncorrelated phase noise
@@
In certain ion traps and superconducting quantum devices, the sequence of $\operatorname{CZ}$ gates is realized with evolution with an Ising-type Hamiltonian:
@@
B.2 Correlated phase noise
@@
  • The unmodified state graph state |$\ket{G}$ with reduced probability coming from $\operatorname{\mathcal{CZ}}$ terms appearing in product of channels defined by Equation 34,
  • graph states with one edge missing as in the case of uncorrelated noise,
@@
$+\frac{p}{2}\sum_{(e,e^{\prime}\in E}^{e\neq e^{\prime}}\sum_{\pm_{1},\pm_{2}}\pm_{1}\pm_{2}\operatorname{\mathcal{CS}}_{e}^{\pm_{1}}\operatorname{\mathcal{CS}}_{e^{\prime}}^{\pm_{2}}[\ket{G}\bra{G}].$
Here, $\lvert E\rvert$ denotes the number of edges in the graph $G$. Note that it is not an ensemble decomposition: in the last sum, some signs are negative. This decomposition is not unique, and in principle different weights for the full, edge-missing, and $\operatorname{\mathcal{CS}}$-modified graph states are possible; however, some of the signs will always be negative, regardless of the choice. However, such decompositions are only low noise approximations for $p\ll\lvert E\rvert^{-1}$.
B.3 Fusion gate noise
@@
Optical realizations [41], while using different mechanisms for entangling, lead to a similar result. If a type 1 entangling gate is used, imperfect mode matching leads affected qubits to be effectively depolarized in the $Z$ basis. This has the exact same result on the output state as local (single-qubit) phase noise, despite different physical origins. A different scheme for generating large cluster states is employed in linear optics experiments. In such systems, the qubits are encoded as photons in different light modes. Two modes are associated with each qubit, and a single photon being in one of them corresponds to the standard computational basis states. Usually, these are polarization (horizontal/vertical) modes with overlapping spatial profiles. Thus, let us denote the creation operators of those modes by ${\hat{h}^{\dagger}}_{i}$ and ${\hat{v}^{\dagger}}_{i}$.
In this context, a Bell pair encoded as the nontrivial two-vertex graph state can be written as
@@
$=\frac{{\hat{h}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}f_{++}+{\hat{v}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}f_{-+}+{\hat{h}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}f_{+-}+{\hat{v}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}f_{--}}{2}\ket{\emptyset},$
where the symbols $f$ denote expressions involving creation operators to arrive at the state |$\ket{\psi}$. The $f_{++}$ prepares the graph state |$\ket{G^{\prime}}$, while $f_{-+}$ creates the state $\prod_{(1,i)\in G}z_{i}\ket{G^{\prime}}$: it is the postselected graph state upon measuring negative sign of the first qubit in the $Z$ basis. The other two operators work similarly; however, $f_{--}$ flips the phase of |$\ket{G^{\prime}}$ only in the qubits that are connected to only $1$ or $2$.
Polarizing beamsplitter exchanges the vertical polarization, keeping the horizontal ones intact ($U{\hat{h}^{\dagger}}_{1,2}={\hat{h}^{\dagger}}_{1,2}U$, $U{\hat{v}^{\dagger}}_{1,2}={\hat{v}^{\dagger}}_{2,1}U$):
$U\ket{G}$ $=\frac{{\hat{h}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}f_{++}+{\hat{v}^{\dagger}}_{2}{\hat{h}^{\dagger}}_{2}f_{-+}+{\hat{h}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{1}f_{+-}+{\hat{v}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}f_{--}}{2}\ket{\emptyset}.$ (42)
@@
$\frac{{\hat{h}^{\dagger}}_{1}f_{++}+{\hat{v}^{\dagger}}_{1}f_{--}}{\sqrt{2}}\ket{\emptyset}.$ (43)
@@
This is the case only if the photons in modes 1 and 2 are indistinguishable after mixing through the polarizing beamsplitter – the procedure relies on the Hong-Ou-Mandel effect. Shall this assumption not be met, the interference required for the fusion to work does not take place, and effectively the two photons originating in mode 1 and 2 are measured independently. If only one photon was observed in the new mode 2, half of the times it arrived from the mode (1,v), while the photon from mode 2 is now localized in the new mode (1,v) and the final state is ${\hat{v}^{\dagger}}_{1}f_{--}\ket{\emptyset}$. The opposite process means that the resulting state ${\hat{h}^{\dagger}}_{1}f_{++}\ket{\emptyset}$.
This effectively creates an ensemble which allows for the following interpretation: the fusion took place as if the photons were indistinguishable, but it was immediately measured in the $Z_{1}$ basis and the result of measurement is not known. This is equivalent to a simple probabilistic $Z_{1}$ flip:
$\overbrace{\Pi_{Z_{1}+}}^{\frac{1}{2}(1+Z_{1})}\rho\Pi_{Z+}+\overbrace{\Pi_{Z_{1}-}}^{\frac{1}{2}({1-Z_{1}})}\rho\Pi_{Z_{1}-}=\frac{\rho+Z_{1}\rho Z_{1}}{2}.$ (44)

### v2

@@ Appendix A Bell pair creation is equivalent to teleportation
Let us assume the following:
  • We fix a graph $G=\{V,E\}$, and choose the two terminal vertices $\{{\operatorname{in}},{\operatorname{out}}\}$ such that $V=\{{\operatorname{in}},{\operatorname{out}}\}\cup I$.
  • The corresponding graph state |$\ket{G}$ undergoes a sequential measurement of $I$ producing a Bell pair across $\{{\operatorname{in}},{\operatorname{out}}\}$. According to the Lemma 1, this means that the correlations between $\{{\operatorname{in}},{\operatorname{out}}\}$ are described by stabilizer operators of |$\ket{G}$ modified by the measurement process (Equation 12).
  • Let $G$ be a part of a larger graph $G^{\prime}$ in the following sense: $G^{\prime}=(V^{\prime},E^{\prime})$ embeds $G$ in such a way that $\{{\operatorname{in}},{\operatorname{out}}\}$ is a separating set. This means that the only connections between $G$ and the rest of $G^{\prime}$ are at these two vertices.
  • Let us denote the vertices in $V^{\prime}\setminus V$ to which both qubits are connected by $\mathcal{N}^{\prime}({\operatorname{in}})$ and $\mathcal{N}^{\prime}({\operatorname{out}})$, respectively, and assume that the two sets are disjoint.
Consider now the following process viewed from the perspective of the graph $G^{\prime}$:
  1. (a)
    the measurement scheme, which within $G$ would produce a Bell pair across the terminal qubits, followed by
  2. (b)
    a product of single qubit unitary operations on the qubits $\{{\operatorname{in}},{\operatorname{out}}\}$, aimed to counteract the effect of measurement and bring back the stabilizer operators into a manageable form, and
  3. (c)
    the measurement of $X_{\operatorname{in}}$.
These operations taken together effectively performs qubit fusion of the terminal qubits. The new state is a graph state where the qubits $V^{\prime}\setminus V$ are intact, and the new neighborhood of the remaining qubit is $\mathcal{N}_{{\operatorname{out}}}\cup\mathcal{N}_{{\operatorname{in}}}$.
@@
If within $G$ a Bell pair is produced across the terminal vertices $\{{\operatorname{in}},{\operatorname{out}}\}$ as a result of the measurement pattern $O=\prod_{i\in I}o_{i}$, this means that there exist three $O$-consistent stabilizers involving the terminal vertices in the sense of Lemma 1. Combinatorical considerations show that one of those must be of a form such that the operators at the terminal vertices are $A_{\operatorname{in}},B_{\operatorname{out}}\in\{X,Y\}$ (but not $Z$), chosen independently.
Thus, there exists an $G$-stabilizer operator of the form ±$\pm A_{\operatorname{in}}B_{\operatorname{out}}\prod_{j\in J}o_{j}=g_{\operatorname{in}}g_{\operatorname{out}}\prod_{j\in J^{*}}g_{j}$, with $J,J^{*}\subset I$. Note that the definition of $g_{i}$ depends on the graph in question (Equation 6): if by $g^{\prime}_{i}$ we denote the generators corresponding to the entire graph $G^{\prime}$, the result is also a stabilizer operator, differing only by the $Z$ operators in the set $\mathcal{N}^{\prime}({\operatorname{in}})\cup\mathcal{N}^{\prime}({\operatorname{out}})$.
Let us now determine the post-measurement stabilizer operator within $G^{\prime}$ by application of Lemma 1. As prescribed by the operator $O$, every qubit $i\in L$ is measured in the basis of $o_{i}$, and the qubit $({\operatorname{in}})$ is measured in the basis of $A_{\operatorname{in}}$. As a result, the post-measurement stabilizer operator is
(±$\left(\pm s_{\operatorname{in}}\prod_{j\in J}s_{j}\right)B_{\operatorname{out}}\mathchoice{}{}{}{}\prod_{j^{\prime}\in\mathcal{N}^{\prime}_{\operatorname{in}}\cup\mathcal{N}^{\prime}_{\operatorname{out}}}\mathchoice{}{}{}{}Z_{j^{\prime}},$ (23)
which can be taken to be a new generator associated with the vertex ${\operatorname{out}}$, and brought to the canonical form of $X_{\operatorname{out}}\prod_{j^{\prime}\in\mathcal{N}^{\prime}_{\operatorname{in}}\cup\mathcal{N}^{\prime}_{\operatorname{out}}}Z_{j^{\prime}}$ by local unitary basis change of the ${\operatorname{out}}$-qubit. The properties of other stabilizers involving the qubits $\{{\operatorname{in}},{\operatorname{out}}\}$ can be proven similarly. They merge to stabilizer operators of only the qubit remaining out-qubit.
Thus, after the measurement within the Bell-generating part $I\cup\{{\operatorname{in}}\}$, the ${\operatorname{out}}$-qubit behaves exactly like it was connected to the neighborhood $\mathcal{N}^{\prime}_{\operatorname{in}}$ as well: the two vertices are fused and the information between them is transferred. ∎
Appendix B Post-measurement correlations and stabilizers compatible with the measurement scheme
Lemma 4.
Fix a graph state |$\ket{G}$ of $N$ qubits and a certain measurement scheme defined in Lemma 1 by a set of qubits $I$ and operators $o_{i}$. The stabilizer operators of |$\ket{G}$ which are compatible with the measurement pattern (that is, the operators share elements with the measurement scheme, see discussion before Equation 12) form a subgroup. This subgroup has $N-\lvert I\rvert$ generators, and therefore fully defines the post-measurement state.
Proof.
@@ Appendix B Post-measurement correlations and stabilizers compatible with the measurement scheme
The stabilizer of |$\ket{G}$, up to the structure of the signs, can be viewed as a linear subspace of $\mathbb{Z}_{2}^{2N}$ – this is known as the binary representation of the stabilizer [7]. The identification is following: if the stabilizer operator is defined as $\prod_{j\in J}g_{j}$, the corresponding element of $\mathbb{Z}_{2}^{2N}$ is $(x_{1},\ldots,x_{N})\oplus(z_{1},\ldots,z_{N})$, with $x_{i}$ being 1 if and only if $i\in J$ and $\vec{z}=A_{G}\vec{x}\mod 2$, where $A_{G}$ is the adjacency matrix of $G$. Thus, if a stabilizer operator contains $X_{i}$, the corresponding numbers are $(x_{i},z_{i})=(1,0)$, for $Z_{i}$ it is $(x_{i},z_{i})=(0,1)$, and $Y_{i}$ corresponds to $(x_{i},z_{i})=(1,1)$; no operator at site $i$ is denoted by $(x_{i},z_{i})=(0,0)$. In this way, the additive algebra of $\mathbb{Z}_{2}^{2}$ mimics the product rules of Pauli operators with only the sign structure missing.
@@
The stabilizer of |$\ket{G}$ corresponds to a special $N$-dimensional subspace in $\mathbb{Z}_{2}^{2N}$ [7]. For any stabilizer operator $Q=\prod_{j\in J}q_{j}$ to be compatible with the measurement scheme $O=\prod_{i\in I}o_{i}$ in the sense of Lemma 1, each operator of $q_{i}$ its Pauli string form with support in $I$ must be $o_{i}$: this introduces constraints to the set of compatible operators. These constraints translated into the language of $\mathbb{Z}_{2}^{2N}$ are linear: if $X_{i}$ is measured at site $i$, the additional constraint is $z_{i}=0$ (this encompasses the cases of $X_{i}$ and identity at site $i$), if $Z_{i}$ is measured, the constraint is $x_{i}=0$, and for $Y_{i}$ it is $x_{i}+z_{i}=0\mod 2$. All of them are independent, and there are $\lvert I\rvert$ of them, one for each measured site. Thus, the linear subspace of $\mathbb{Z}_{2}^{2N}$ corresponding to measurement-compatible stabilizers has dimension $N-\lvert I\rvert$; any basis of this subspace corresponds to generators of the post-measurement state stabilizer. ∎
@@
B.1 Discussion of measurement outcomes
Figure 10: Some measurement patterns have a restricted set of accepted outcomes. If we measure the pairs of qubits associated to the vertices in the gray areas in Pauli $X$ basis, the products of the outcomes have to be equal to one. This indicates which combination of signs are allowed: The accepted signs are shown in the green areas while the rejected signs are shown in the red areas. We postselect by discarding all states whose measurement outcomes have signs which disagree with the target graph state.
[…]
In Sections 5 and 6, we perform measurements on noisy graph states and postselect on measurement outcomes. That is, we consider all stabilizer operators with support within the measurement pattern and check the outcome for consistency. If the outcomes do not match the expectation, we discard the resulting state. This concept is also pictured in Figure 2.
B.2 Patterns for specific graph families
@@
Both cases are consistent with a maximally entangled two-qubit state being produced in the terminal qubits: for the odd-$n$ case, the existence of stabilizers described by LABEL:eqn:termoddpathcorr ensures that further measurement of $X$, $Y$ or $Z$ at qubit 00 is perfectly (anti)correlated with the outcome of measurement of the same operator at qubit $n+1$.
@@
This class of graphs includes the path graph (each layer has only a single qubit), the crazy graph (the layers 00 and $n+1$ contain one qubit each, every other has two) and twisted graphs (the qubit count alternates between 1 and 2) as well. Suppose that all the qubits in every internal ($i=1,\ldots,n$) layer are measured in the $X$ basis. The stabilizer operators that determine the remaining correlations must be consistent with this structure. A particularly simple set of such operators exists: first, observe that if a layer $k$ contains at least two qubits $i$ and $i^{\prime}$, then $X_{k,i}X_{k,i^{\prime}}$ is a stabilizer operator. Thus, the signs of outcomes have to be equal within a layer. Therefore, for any choice of inner-layer qubit index $(i_{k})$ the direct analogue of LABEL:eqn:stabextended holds:
@@
are stabilizer operators. For symmetry, we assume the twisted graph structure exists only for an odd number of layers, so that the terminal ones consist only of one qubit each. Thus, the terminal layers 00 and $n+1$ always consist only of one qubit each, and the inner indices are omitted. Consequently, if $s_{k}$ in LABEL:eqn:termoddpathcorr are interpreted as signs of outcomes at layer $k$ (which are all equal within a layer due to the structure of the stabilizer operators mentioned previously), the same post-measurement stabilizer structure appears.
In the case of the crazy graph, an additional observation can be made: loss of any single internal (not adjacent to the terminal qubits) edge yields the same post-measurement stabilizer structure, since the indices $(i_{k})$ in LABEL:eqn:stabextendedgeneral can be chosen arbitrarily. Thus, the crazy graph family is especially tolerant to single edge loss: it does not affect the post-measurement state, if the outcomes are postselected on the ideal sign structure. This can be generalized to the loss of any subset of edges not adjacent to each other and the terminal qubits.
Appendix C Derivation for the noise models
C.1 Uncorrelated phase noise
@@
In certain ion traps and superconducting quantum devices, the sequence of $\operatorname{CZ}$ gates is realized with evolution with an Ising-type Hamiltonian [41, 42, 43, 44] :
@@
C.2 Correlated phase noise
@@
  • The unmodified state graph state |$\ket{G}$, with reduced probability coming from $\operatorname{\mathcal{CZ}}$ terms appearing in product of channels defined by Equation 31,
  • graph states with one edge missing, as in the case of uncorrelated noise,
@@
$+\frac{p}{2}\sum_{(e,e^{\prime}\in E}^{e\neq e^{\prime}}\sum_{s_{1},s_{2}=\pm 1}s_{1}s_{2}\operatorname{\mathcal{CS}}_{e}^{s_{1}}\operatorname{\mathcal{CS}}_{e^{\prime}}^{s_{2}}[\ket{G}\bra{G}].$
Here, $\lvert E\rvert$ denotes the number of edges in the graph $G$. Note that this decomposition is not convex and cannot be interpreted as an ensemble: the minus signs in front of the decomposition terms prevent this. Signs appear as a result of discrete decomposition of controlled-phase and subsequent averaging; see Appendix C for details. In this decomposition, $G-e\coloneqq(V,E\setminus\{e\})$ is the graph with edge $e$ removed, and ${\operatorname{\mathcal{CS}}^{\pm}}$ is the application of one of the two unitary square roots of $\operatorname{CZ}$:
where
$\sqrt{\operatorname{CZ}^{\pm}}\coloneqq\frac{1}{2}\left[\left(1+e^{\pm i\frac{\pi}{2}}\right)\operatorname{\mathds{1}}+\left(1-e^{\pm i\frac{\pi}{2}}\right)\operatorname{CZ}\right]=\begin{pmatrix}1&&&\\ &1&&\\ &&1&\\ &&&\pm i\end{pmatrix}.$ (37)
This decomposition is not unique, and in principle different weights for the full, edge-missing, and $\operatorname{\mathcal{CS}}$-modified graph states are possible; however, some of the signs will always be negative, regardless of the choice. All such decompositions are only low noise approximations for $p\ll\lvert E\rvert^{-1}$.
C.3 Fusion gate noise
@@
A different scheme for generating large cluster states is employed in linear optics experiments [56]. In such systems, the qubits are encoded as photons in different light modes. Two modes are associated with each qubit, and a single photon being in one of them corresponds to the standard computational basis states. Then, the graph states are built by qubit fusion of initially generated Bell pairs. If a type 1 entangling gate is used for this, imperfect mode matching leads affected qubits to be effectively depolarized in the $Z$ basis. This has the exact same result on the output state as local (single-qubit) phase noise, despite different physical origins.
Usually, the two modes defining qubit degrees of freedom are polarization (horizontal/vertical) modes with overlapping spatial profiles. Thus, let us denote the creation operators of those modes by ${\hat{h}^{\dagger}}_{i}$ and ${\hat{v}^{\dagger}}_{i}$. In this context, a Bell pair encoded as the nontrivial two-vertex graph state can be written as
@@
$=\frac{{\hat{h}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}\hat{f}_{++}+{\hat{v}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}\hat{f}_{-+}+{\hat{h}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}\hat{f}_{+-}+{\hat{v}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}\hat{f}_{--}}{2}\ket{\emptyset},$
where the symbols $\hat{f}$ denote expressions involving creation operators to arrive at the state |$\ket{\psi}$. The $\hat{f}_{++}$ prepares the graph state |$\ket{G^{\prime}}$, while $\hat{f}_{-+}$ creates the state $\prod_{(1,i)\in G}z_{i}\ket{G^{\prime}}$: it is the postselected graph state upon measuring negative sign of the first qubit in the $Z$ basis. The other two operators work similarly; however, $\hat{f}_{--}$ flips the phase of |$\ket{G^{\prime}}$ only in the qubits that are connected to only $1$ or $2$.
Polarizing beamsplitter, described mathematically by a unitary operator $U$, exchanges the vertical polarization, keeping the horizontal ones intact ($U{\hat{h}^{\dagger}}_{i}={\hat{h}^{\dagger}}_{i}U$, $U{\hat{v}^{\dagger}}_{1}={\hat{v}^{\dagger}}_{2}U$, and $U{\hat{v}^{\dagger}}_{2}={\hat{v}^{\dagger}}_{1}U$). First part of the fusion process leads to:
$U\ket{G}$ $=\frac{{\hat{h}^{\dagger}}_{1}{\hat{h}^{\dagger}}_{2}\hat{f}_{++}+{\hat{v}^{\dagger}}_{2}{\hat{h}^{\dagger}}_{2}\hat{f}_{-+}+{\hat{h}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{1}\hat{f}_{+-}+{\hat{v}^{\dagger}}_{1}{\hat{v}^{\dagger}}_{2}\hat{f}_{--}}{2}\ket{\emptyset}.$ (41)
@@
$\frac{{\hat{h}^{\dagger}}_{1}\hat{f}_{++}+{\hat{v}^{\dagger}}_{1}\hat{f}_{--}}{\sqrt{2}}\ket{\emptyset}.$ (42)
@@
This is the case only if the photons in modes 1 and 2 are indistinguishable after mixing through the polarizing beamsplitter – the procedure relies on the Hong-Ou-Mandel effect. Shall this assumption not be met, the interference required for the fusion to work does not take place, and effectively the two photons originating in mode 1 and 2 are measured independently. If only one photon was observed in the new mode 2, half of the times it arrived from the mode (1,v), while the photon from mode 2 is now localized in the new mode (1,v) and the final state is ${\hat{v}^{\dagger}}_{1}\hat{f}_{--}\ket{\emptyset}$. The opposite process means that the resulting state ${\hat{h}^{\dagger}}_{1}\hat{f}_{++}\ket{\emptyset}$.
This effectively creates an ensemble which allows for the following interpretation: the fusion took place as if the photons were indistinguishable, but it was immediately measured in the $Z_{1}$ basis and the result of measurement is not known. This is equivalent to a simple probabilistic $Z_{1}$ flip, described by the following channel:
$\rho\mapsto\overbrace{\Pi_{Z_{1}+}}^{\frac{1}{2}(1+Z_{1})}\rho\Pi_{Z+}+\overbrace{\Pi_{Z_{1}-}}^{\frac{1}{2}({1-Z_{1}})}\rho\Pi_{Z_{1}-}=\frac{\rho+Z_{1}\rho Z_{1}}{2}.$ (43)
@@
Appendix D Cancellation of correlated edge noise effects in twist and crazy graph families
If for the square graph presented in Figure 2a, instead of the perfect controlled-$Z$ unitary, the controlled phase
$\operatorname{C\Phi}(\pi+\varepsilon)=\operatorname{diag}(1,1,1,-1\exp(i\varepsilon))$ (44)
is used with the same $\varepsilon$ for all edges, the result is a four-qubit weighted graph state which we denote by |$\ket{G}_{\varepsilon}$. If qubits 1 and 3 are measured in the $X$ basis, we postselect on the measurement outcomes which are equal. With the weighted graph state |$\ket{G}_{\varepsilon}$ as the initial state, let us denote the unnormalized post-measurement state by
|$\ket{\xi}_{\pm}=\frac{\mathbbm{1}\pm X_{1}}{2}\frac{\mathbbm{1}\pm X_{3}}{2}\ket{G}_{\varepsilon}.$ (45)
Note the equality of signs $\pm$ appearing in two projection operators: we postselect on observing this type of outcome, since unequal signs can not appear in the ideal ($\varepsilon=0$) case. The relevant expectation values defining the end correlations (with $\langle O\rangle_{\pm}$ denoting the expectation value of $O$ for the unnormalized state |$\ket{\xi}_{\pm}$) across qubits $2$ and $4$ are the following, in the lowest nontrivial noise order:
$\langle Z_{2}Z_{4}\rangle_{+}$ $=1-\varepsilon^{2}+O(\varepsilon^{4}),$ $\langle Z_{2}Z_{4}\rangle_{-}$ $=-1+\frac{\varepsilon^{2}}{2}+O(\varepsilon^{4}),$
$\langle X_{2}X_{4}\rangle_{+}$ $=1-3\varepsilon^{2}+O(\varepsilon^{4}),$ $\langle X_{2}X_{4}\rangle_{-}$ $=-1+\frac{\varepsilon^{2}}{2}+O(\varepsilon^{4}),$
$\langle Y_{2}Y_{4}\rangle_{+}$ $=-1+3\varepsilon^{2}+O(\varepsilon^{4}),$ $\langle Y_{2}Y_{4}\rangle_{-}$ $=1-\frac{\varepsilon^{2}}{2}+O(\varepsilon^{4}),$
$\||\xi\rangle_{+}\|^{2}$ $=1-\varepsilon^{2}+O(\varepsilon^{4}),$ $\||\xi\rangle_{-}\|^{2}$ $=1-\frac{\varepsilon^{2}}{2}+O(\varepsilon^{4}).$ (46)
Note that for the |$\ket{\xi}_{-}$ outcome the squared length $\||\xi\rangle_{-}\|^{2}$ is proportional to the unnormalized expectation values: thus, for the physical correlations the dominant noise effects are canceled and only higher order terms ($O(\varepsilon^{4})$) remain. On the other hand, if the outcome is |$\ket{\xi}_{+}$, the noise is amplified compared to postselection on only stabilizer-consistent outcomes. Thus, further postselection can amplify the entanglement quality of the remaining qubits at the cost of discarding certain outcomes and reduced production rates.
The double twist graph can be viewed as multiple such squares stacked together by the terminal vertices: thus, a proper postselection can mitigate the first nontrivial noise effects completely (see Figure 7). A similar structure appears for the crazy graph: it is much more complex to analyze it, but computer algebra systems provide evidence of a periodic structure (also visible Figure 7) in of susceptibilities as the length of the graph increases.